OpenStudy (anonymous):

3.6q-2.6-3.2q=0.6q-2.4

OpenStudy (theslytherinhelper):

Did you try combining like terms?

OpenStudy (anonymous):

yes i did but i think i did it wrong

OpenStudy (theslytherinhelper):

Why? Let's try it. 3.6q-2.6-3.2q=0.6q-2.4 .4q - 2.6 = 0.6q - 2.4 -2.6 = [1]q - 2.4 -5 = q Is that what you got? If not, then it's the right answer.

OpenStudy (anonymous):

where di you get the 4q from? when i did it, i think i solved it wrong

OpenStudy (theslytherinhelper):

The .4q is from subtracting 3.6 - 3.2 q

OpenStudy (anonymous):

*did

OpenStudy (theslytherinhelper):

You're basically combining like terms the entire way through...

OpenStudy (anonymous):

oh ok cause i did that but it didnt come out that way

OpenStudy (theslytherinhelper):

Did you find your error?

OpenStudy (anonymous):

yeah i did thank you